# IMUFace: Real-Time, Low-Power, Continuous 3D Facial Reconstruction Through Earphones

### v1

@@ Abstract
The potential of facial expression reconstruction technology is significant, with applications in various fields such as human-computer interaction, affective computing, and virtual reality. Recent studies have proposed using ear-worn devices for facial expression reconstruction to address the environmental limitations and privacy concerns associated with traditional camera-based methods. However, these approaches still require improvements in terms of aesthetics and power consumption. This paper introduces a system called IMUFace. It uses inertial measurement units (IMUs) embedded in wireless earphones to detect subtle ear movements caused by facial muscle activities, allowing for covert and low-power facial reconstruction. A user study involving 12 participants was conducted, and a deep learning model named IMUTwinTrans was proposed. The results show that IMUFace can accurately predict users’ facial landmarks with a precision of 2.21 mm, using only five minutes of training data. The predicted landmarks can be utilized to reconstruct a three-dimensional facial model. IMUFace operates at a sampling rate of 30 Hz with a relatively low power consumption of 58 mW. The findings presented in this study demonstrate the real-world applicability of IMUFace and highlight potential directions for further research to facilitate its practical adoption.
Index Terms:
@@
I Introduction
@@
Facial expressions are essential in interpersonal communication because they contain various information, including emotional states [1, 2] and health conditions [3, 4]. Consequently, facial expression reconstruction is widely applied in areas requiring facial analysis. It has become a fundamental technology in fields such as human-computer interaction (HCI), virtual reality (VR), affective computing, and healthcare. For example, facial expression reconstruction can enhance users’ real-time perception of facial expressions and emotions in VR, providing a more immersive experience [5]. In metaverse applications, it can rapidly establish understanding and connections within virtual spaces [6, 7]. Additionally, it can enable silent speech interfaces, facilitating seamless human-computer interaction [8].
Early facial expression reconstruction methods typically relied on manual annotation and geometric models. These traditional methods had low automation and were complex to implement, thus facing significant limitations. With advancements in computer vision and machine learning, automated facial expression reconstruction advances the field. Traditional vision-based methods [9, 10, 11] can accurately locate facial landmarks and generate high-quality facial animations. However, the use of direct camera angles and environmental factors imposes constraints on their efficacy. In order to ensure accuracy, cameras must often face users directly, which limits the applicability of these methods in situations involving varying lighting, resolution, or occlusion conditions. Additionally, the direct capture of users’ faces may give rise to privacy concerns, particularly in contexts where 3D head movements and dynamic environments are involved, as such visual approaches are less suitable in such scenarios.
These limitations have driven a shift toward non-visual facial expression reconstruction methods. Researchers have proposed combining wearable devices and multimodal sensors to enable more flexible and privacy-friendly solutions, such as necklaces [12], hats [13], glasses [14, 15, 16] and ear-mounted devices [17, 18, 19, 20, 21]. Earlier works, like C-Face [22], used miniature ear-mounted cameras to capture facial expressions, offering flexible deployment options. Nonetheless, privacy concerns persist due to camera mobility and proximity to others. Subsequently, BioFace-3D [19] and EarIO [20] introduced non-visual 3D facial reconstruction through ear-mounted devices. However, their designs may cause discomfort by protruding behind the head or slightly touching the face. Later, EARFace [21] embedded acoustic sensors within earplug designs, yielding higher usability scores. However, high power consumption limits acoustic sensors’ viability for prolonged use. Therefore, developing a portable, unobtrusive, low-power system for real-time, continuous facial expression reconstruction remains challenging.
We developed a deep learning model called IMUTwinTrans to achieve 3D facial reconstruction beyond the limitations of cameras. This model, based on the ConvTransformer architecture, is capable of simultaneously leveraging the temporal and frequency features of IMU signals to establish the correspondence between IMU signals and facial landmarks. Additionally, the model adopts a lightweight design to meet the requirements for real-time performance. During testing, the well-trained model can localize facial landmarks directly from IMU signals without visual input. The predicted facial landmarks are then filtered and fitted into a generic 3D head model to generate continuous 3D facial animation. Fig. 1 shows examples of 3D facial reconstruction using the proposed IMUFace system, demonstrating its ability to monitor various parts of the face across different facial motions.
@@
II Related Work
In non-wearable 3D face reconstruction, various algorithms have been developed to address challenges posed by unconstrained photo collections and achieve high-fidelity reconstructions. These methods can be categorized based on their core algorithmic approaches: morphable models, photometric stereo techniques, and deep learning-based solutions.
Morphable models (3DMMs) have been a cornerstone in 3D face reconstruction due to their ability to represent facial geometry and texture in a compact form. Roth et al. [23] introduced an adaptive method for reconstructing 3D face surfaces from unconstrained photo collections using a 3DMM. They employed a coarse-to-fine fitting strategy that allowed the model to adapt even to low-quality image sets, thereby expanding its applicability. This approach improved the quality of reconstructions and handled variations in pose, expression, and illumination more effectively. Roth et al. further advanced this work in [24], where they proposed an algorithm capable of generating accurate 3D face models from diverse 2D images without metadata about cameras or timing. By leveraging a standard field-based Laplace editing technique, they ensured robustness across varied imaging conditions.
Photometric stereo is another key technique used in 3D face reconstruction, especially when dealing with extensive photo collections. Roth et al. [24] utilized this method to capture detailed surface normals from multiple viewpoints, enabling the construction of accurate 3D models. The integration of photometric stereo with landmark constraints provided a powerful means to drive surface reconstruction while maintaining fidelity to the input images. Moreover, Xie et al. [25] explored photometric stereo in the context of smartphone-based reconstruction, demonstrating how calibrated screen lighting can be used to recover fine-grained textures and micro facial features. Their method addresses practical limitations such as hand jitter through the use of face landmarks for alignment and outlier normal elimination techniques.
@@ II Related Work
Deep learning has revolutionized 3D face reconstruction by enabling high-fidelity outputs from single images. Gecer et al. [26] introduced GANFIT, which combines Generative Adversarial Networks (GANs) and Deep Convolutional Neural Networks (DCNNs) to reconstruct facial texture and shape. GANFIT trains a powerful generator of facial texture in UV space and optimizes the parameters under the supervision of pretrained deep identity features. This approach achieves photorealistic and identity-preserving reconstructions with unprecedented detail. Built upon GANFIT, Gecer et al. [27] developed Fast-GANFIT. a faster and more robust version introducing a self-supervised regression-based initialization scheme to expedite the fitting process and enhance stability.
These studies collectively illustrate the evolution of non-wearable 3D face reconstruction algorithms, each offering unique strengths. The versatility of morphable models, the precision of photometric stereo techniques, and the fidelity of deep learning-based solutions all contribute to this advancement. This variety in methodologies highlights the continuous innovation and adaptability present in the field.
II-A 3D Face Reconstruction with Different Models
The 3D Morphable Model (3DMM) [28] serves as the foundational basis for 3D face reconstruction. It represents facial geometry through a linear combination of orthogonal basis functions, with a fixed number of points to depict a face. However, traditional 3DMM approaches require optimizing to solve for relevant coefficients, which is often complex and computationally intensive. To address these challenges, increasing attention has been directed towards leveraging deep learning techniques to directly regress these coefficients, enabling more efficient and precise 3D face reconstruction.
@@
In the realm of deep learning, CNNs have been extensively applied to 3D face reconstruction. For instance, the paper Richardson et al. [29]presents an end-to-end CNN framework that reconstructs detailed facial geometry in a coarse-to-fine manner. Similarly, Tian et al. [30] introduce a CNN-based method for reconstructing 3D faces from two facial images, demonstrating superior performance across various datasets.
@@
Beyond CNNs, Encoder-Decoder architectures have also been employed in 3D face reconstruction. Ji et al. [31] proposes a multi-task encoder-decoder framework based on a Siamese network to learn view-consistent shapes and textures from a single image. Meanwhile, Li et al. [32] enhanced reconstruction accuracy by using a dual-pathway encoder-decoder network to distinguish between different attributes.
@@
Recently, Transformer models have made significant strides in 3D face reconstruction. Chen et al. [33] utilizes a conditional Generative Adversarial Network (cGAN) and a novel mesh transformer for cross-domain face synthesis and reconstruction. It introduces a reprojection consistency loss for self-supervised learning. Another study, Yaermaimaiti et al. [34] integrates ResNet and Transformer modules, leveraging self-attention mechanisms to improve parameter prediction accuracy and significantly enhance the reconstruction results.
II-B Non-wearable 3D Face Reconstruction
The research field of 3D face reconstruction employs various data sources generated by different devices to improve reconstruction quality and applicability. Traditional 3D face reconstruction methods primarily rely on external imaging devices, such as RGB cameras, utilizing extensive photo collections for reconstruction. For instance, Roth et al. [23] present a method capable of reconstructing a 3D face model of an individual from photo collections captured under varied poses and lighting conditions while also extracting albedo information. This approach achieves superior experimental results by fitting a 3D Morphable Model to form a personalized template and developing a novel photometric stereo formulation. Similarly, Roth et al. [24] introduce an algorithm that generates an accurate 3D face model from unconstrained collections of 2D images. This algorithm employs a photometric stereo method along with a novel normal field-based Laplace editing technique, leveraging images captured in diverse poses, and demonstrates both qualitative and quantitative effectiveness.
II-C Wearable 3D Face Reconstruction
In recent years, researchers have begun to explore 3D face reconstruction using wearable miniature cameras, aiming to facilitate face reconstruction in real-life scenarios while minimizing user interference. Chen et al. [22] propose an ear-mounted wearable device that captures subtle changes in facial contours using two miniature cameras and utilizes a deep learning model to continuously output facial feature points, enabling continuous facial expression reconstruction. Additionally, Aoki et al. [35] introduces a glasses-shaped device that captures the user’s facial expressions through two miniature cameras and three mirrors, allowing for the recording and recognition of facial expressions without obstructing daily activities. The system demonstrated high accuracy in emotion recognition, underscoring its applicability to facial image indexing. These studies highlight the expanding potential of wearable devices in 3D face reconstruction, opening new possibilities for practical applications.
The transition from visual methods to multi-modal approaches has led to innovative techniques for monitoring facial expressions and movements using sensors like proximity sensors, EMG, pressure sensors, EOG, and audio signals. Notably, smart glasses and earphones utilizing IMU and acoustic sensors have emerged as effective tools for 3D reconstruction. Wu et al. [19] uses a single-earpiece biosensing system to track facial landmarks and render 3D animations from biosignals without visual input, achieving accuracy comparable to camera-based solutions. Li et al. [16] employ IMUs on smart glasses to estimate action unit intensities for real-time facial reconstruction with low power consumption. Zhang et al. [21] leverages in-ear acoustic sensors to track facial landmarks by detecting changes in the ear canal shape, providing robust and privacy-preserving 3D facial reconstruction. These advancements demonstrate the potential of multi-modal sensors in offering accurate, unobtrusive, and privacy-sensitive facial tracking suitable for diverse applications.
III Background
In this section, we introduce the background information on the following aspects: (1) facial motion and the muscles involved; (2) principle of operation of IMUFace.
III-A Facial Muscles and Expressions
Facial expressions are a crucial element of human communication, providing insight into emotions and intentions that others can perceive and interpret. The human face, equipped with 43 distinct muscles, is capable of producing more than 10,000 different expressions. These muscles, known as mimetic muscles, are organized into three major categories: orbital, nasal, and oral. The orbital muscles control eyelid movements, the nasal group manages movements related to the nose, and the oral muscles facilitate motions of the lips and mouth. Through the interaction of these muscle groups, the human face can depict a wide array of expressions, such as happiness, sadness, surprise, and fear.
The complexity of facial expressions arises from the coordinated activity of these muscle groups, which pull on the facial skin to create the visual cues associated with emotions. For instance, the frontalis muscle raises the eyebrows in surprise, while the temporalis and lateral pterygoid muscles contribute to jaw movements in expressions of amazement or joy.
@@ III-A Facial Muscles and Expressions
III-B Sensing Principle
The relationship between facial expressions and the movement of facial muscles provides a foundation for innovative sensing technologies that capture these expressions indirectly. When facial muscles contract, they alter visible expressions and induce subtle changes in related anatomical structures, such as the ear canal, due to interconnected tissue and bone dynamics.
@@ III-B Sensing Principle
For example, jaw movements involved in expressions like surprise or joy can affect the mandibular condyle, leading to changes in the shape of the ear canal. Similarly, the temporalis muscle, a major muscle in the head, transmits mechanical artifacts from facial expressions toward the ear area.
@@
An Inertial Measurement Unit (IMU) embedded in an ear-mounted device, like a wireless earbud, can be employed to capture these changes. As facial expressions lead to deformations of the ear canal, the earbud’s position and orientation are altered. The IMU detects these variations through accelerometers and gyroscopes, translating these subtle physical changes into measurable data. Each facial expression has a unique motion pattern, differentiating distinct facial expressions based on the IMU’s output signals. By employing advanced machine learning algorithms, these signals can be analyzed to identify specific facial expressions, offering a practical method to decode them via ear-mounted sensors. This approach opens new possibilities for unobtrusive facial expression recognition, leveraging the natural connection between facial musculature and ear canal dynamics.

### v3

# IMUFace: Towards Always-On 3D Facial Reconstruction via Earphone Inertial Sensing

@@ Abstract
Facial expression reconstruction technology offers considerable potential in areas like human-computer interaction, affective computing, and virtual reality. To tackle the privacy challenges and environmental constraints inherent in traditional camera-based systems, researchers have recently introduced ear-worn devices as a viable solution. Nevertheless, these methods still demand enhancements, particularly in aspects such as design appeal and energy efficiency, to achieve broader applicability and practical use. This paper introduces a system called IMUFace. It uses inertial measurement units (IMUs) embedded in wireless earphones to detect subtle ear movements caused by facial muscle activities, allowing for covert and low-power facial reconstruction. A user study involving 12 participants was conducted, and a deep learning model named IMUTwinTrans was proposed. The results show that IMUFace can accurately predict users’ facial landmarks with a precision of 2.21 mm, using only five minutes of training data. The predicted landmarks can be utilized to reconstruct a three-dimensional facial model. IMUFace operates at a sampling rate of 30 Hz with a relatively low power consumption of 58 mW. The findings presented in this study demonstrate the real-world applicability of IMUFace and highlight potential directions for further research to facilitate its practical adoption.
keywords:
@@
1 Introduction
Facial expressions play a crucial role in interpersonal communication, as they convey a wide range of information, including emotional states [bassili1979emotion, krumhuber2023role] and physical health conditions [argaud2018facial, roter2006expression]. As a result, facial expression reconstruction has become an essential tool for applications requiring facial analysis. This technology is fundamental to domains such as human-computer interaction (HCI), virtual reality (VR), affective computing, and healthcare. For instance, in VR, it improves users’ real-time perception of emotions and facial expressions, creating a more immersive experience [kumari2015facial]. In metaverse platforms, it enables faster understanding and connection within virtual environments [cha2020real, frith2009role]. It also supports silent speech interfaces, allowing seamless communication between humans and computers [denby2010silent].
Traditional methods for reconstructing facial expressions often relied on manual annotation or geometric modeling. These early approaches were labor-intensive, difficult to automate, and prone to practical limitations. However, with advancements in machine learning and computer vision, automated methods have significantly improved the field. Vision-based techniques [cootes2001active, cootes1995active, wang2020deep] enable precise facial landmark detection and high-quality animation generation. Nevertheless, their reliance on direct camera views introduces several challenges. These methods require users to face the camera under controlled conditions, making them less effective in environments with varying lighting, resolution, or occlusion. Additionally, the direct capture of facial imagery raises privacy concerns, particularly in dynamic or sensitive scenarios involving 3D head movements.
To address these drawbacks, researchers have shifted toward non-visual approaches for facial expression reconstruction. Wearable devices combined with multimodal sensors have emerged as a promising solution, offering greater flexibility and enhanced privacy. Examples include necklaces [chen2021neckface], hats [bello2023inmyface], glasses [xie2021acoustic, li2024eyeecho, li2024auglasses], and ear-mounted devices [choi2022ppgface, song2022facelistener, wu2021bioface, li2022eario, zhang2023earphone]. Earlier designs, such as C-Face [chen2020c], utilized miniature cameras mounted on the ear to capture expressions, providing a flexible deployment option. However, privacy concerns persisted due to the camera’s proximity and mobility. Subsequent innovations like BioFace-3D [wu2021bioface] and EarIO [li2022eario] introduced non-visual 3D facial reconstruction with ear-mounted devices, though their designs sometimes caused discomfort due to protrusions or facial contact. More recent developments, such as EARFace [zhang2023earphone], incorporated acoustic sensors into earplug designs, achieving higher usability ratings. Nonetheless, the high power consumption of acoustic sensors remains a limitation for extended use. Therefore, creating a portable, comfortable, and energy-efficient system capable of real-time facial expression reconstruction continues to be a significant challenge.
@@ 1 Introduction
We developed a deep learning model called IMUTwinTrans to achieve 3D facial reconstruction beyond the limitations of cameras. This model, based on the ConvTransformer architecture, is capable of simultaneously leveraging the temporal and frequency features of IMU signals to establish the correspondence between IMU signals and facial landmarks. Additionally, the model adopts a lightweight design to meet the requirements for real-time performance. During testing, the well-trained model can localize facial landmarks directly from IMU signals without visual input. The predicted facial landmarks are then filtered and fitted into a generic 3D head model to generate continuous 3D facial animation. Fig. 1 showcases 3D facial reconstruction by the IMUFace system and, together with the real-time IMU gyroscope signals (X, Y, Z axes), highlights how subtle angular changes from facial muscle movements produce distinct peaks and troughs that correspond to different expressions, confirming both the sensor’s high sensitivity and the physical meaning of the IMU data for deep learning–based facial reconstruction.
@@
2 Related Work
The field of non-wearable 3D face reconstruction has advanced significantly, driven by developments in morphable models, photometric stereo methods, and deep learning approaches. These methods address challenges such as variations in pose, illumination, and image quality, enabling high-fidelity reconstruction from unconstrained inputs.
Morphable Models serve as a foundational tool for compactly representing facial geometry and texture. Roth et al. [roth2016adaptive] introduced an adaptive 3DMM-based method for reconstructing faces from unconstrained photo collections. Their approach utilized a coarse-to-fine fitting strategy, enhancing robustness against low-quality images and diverse imaging conditions. An earlier work [roth2015unconstrained] proposed an algorithm for generating accurate 3D models from 2D images without relying on metadata, incorporating Laplace editing to improve resilience to pose and illumination variations.
Photometric Stereo techniques have been pivotal in capturing surface normals from multiple viewpoints, contributing to detailed facial reconstructions. For instance, Roth et al. [roth2015unconstrained] integrated photometric stereo with landmark constraints to achieve higher reconstruction accuracy. Xie et al. [xie2021scifi] demonstrated a practical smartphone-based method using calibrated screen lighting, addressing challenges like hand jitter through alignment and outlier removal techniques.
@@ 2 Related Work
Deep Learning has transformed 3D face reconstruction, enabling high-fidelity results from single images. Gecer et al. [gecer2019ganfit] combined Generative Adversarial Networks (GANs) with Deep Convolutional Neural Networks (DCNNs) in GANFIT, achieving photorealistic reconstructions of facial shape and texture. Fast-GANFIT [gecer2021fast], their subsequent work, introduced a self-supervised regression-based initialization scheme, improving speed and stability.
2.1 3D Face Reconstruction with Different Models
The 3D Morphable Model (3DMM) [blanz2023morphable] remains a cornerstone in 3D face reconstruction. While traditional methods optimize coefficients for reconstruction, which is computationally expensive, recent advancements utilize deep learning to directly predict these coefficients, offering greater efficiency and precision.
@@
Convolutional Neural Networks (CNNs) are widely employed in this field. Richardson et al. [richardson2017learning] developed an end-to-end CNN framework for reconstructing facial geometry in a coarse-to-fine manner. Tian et al. [tian2018landmark] presented a CNN-based method for reconstructing 3D faces from two images, demonstrating superior performance on benchmark datasets.
@@
Encoder-Decoder Architectures have also gained prominence. Ji et al. [ji2020view] proposed a multi-task encoder-decoder framework based on a Siamese network to ensure view-consistent reconstruction of facial shapes and textures from single images. Similarly, Li et al. [li2020novel] enhanced accuracy by employing a dual-pathway encoder-decoder network that separates facial attributes.
@@
Transformer Models have recently achieved remarkable progress. Chen et al. [chen2022transformer] utilized a conditional Generative Adversarial Network (cGAN) combined with a mesh transformer for cross-domain face synthesis, introducing a reprojection consistency loss for self-supervised learning. Yaermaimaiti et al. [yaermaimaiti2024research] integrated ResNet and Transformer modules, leveraging self-attention mechanisms to improve parameter prediction and reconstruction quality.
2.2 Non-wearable 3D Face Reconstruction
Traditional methods in 3D face reconstruction rely on external imaging devices like RGB cameras to process extensive photo collections. Roth et al. [roth2016adaptive] proposed a method for reconstructing personalized 3D face models from photos captured under varying poses and lighting conditions. Their approach employed a 3D Morphable Model and a novel photometric stereo formulation to achieve high-quality results. Furthermore, Roth et al. [roth2015unconstrained] introduced an algorithm that integrates photometric stereo with Laplace editing to handle unconstrained 2D image collections effectively.
2.3 Wearable 3D Face Reconstruction
Recent research has focused on wearable devices for 3D face reconstruction, enabling real-time applications with minimal interference. Chen et al. [chen2020c] designed an ear-mounted device equipped with two miniature cameras to capture subtle facial contour changes. Their deep learning model continuously outputs facial feature points, enabling real-time expression reconstruction. Aoki et al. [aoki2021facerecglasses] developed a glasses-shaped device featuring two cameras and three mirrors to capture facial expressions unobtrusively, achieving high accuracy in emotion recognition.
Multimodal Approaches have further enhanced wearable devices. Wu et al. [wu2021bioface] introduced a single-earpiece biosensing system that tracks facial landmarks and generates 3D animations from biosignals, achieving performance comparable to camera-based methods. Li et al. [li2024auglasses] utilized IMUs in smart glasses to estimate action unit intensities, enabling real-time, low-power facial reconstruction. Zhang et al. [zhang2023earphone] employed in-ear acoustic sensors to track facial landmarks by detecting ear canal shape changes, providing a privacy-preserving solution for 3D facial reconstruction.
3 Background
This section provides an overview of the foundational concepts, covering: (1) the role of facial muscles in generating expressions and (2) the operational principle of IMUFace.
@@ 3 Background
3.1 Facial Muscles and Expressions
Facial expressions play an essential role in human communication by conveying emotions and intentions that can be observed and interpreted by others. The human face contains 43 distinct mimetic muscles, which collectively enable the production of over 10,000 unique expressions. These muscles are categorized into three primary groups: orbital, nasal, and oral. The orbital muscles are responsible for eyelid movements, the nasal muscles manage actions around the nose, and the oral muscles control the lips and mouth. The coordination of these groups enables the depiction of a broad range of emotions, such as happiness, sadness, surprise, and fear.
The intricate nature of facial expressions stems from the synchronized activity of these muscle groups, which exert forces on the skin to form visible emotional cues. For instance, the frontalis muscle elevates the eyebrows during expressions of surprise, while jaw movements in amazement or joy involve the temporalis and lateral pterygoid muscles.
3.2 Sensing Principle
The interplay between facial muscle activity and associated anatomical changes provides the basis for innovative sensing methods that indirectly capture facial expressions. Muscle contractions during expressions not only generate visible changes but also induce subtle structural alterations in nearby areas, such as the ear canal, due to the dynamics of interconnected tissues and bones.
@@ 3.2 Sensing Principle
For example, jaw movements that occur during expressions like joy or surprise influence the mandibular condyle, causing changes in the ear canal’s shape. Similarly, the temporalis muscle, a key component of the head’s musculature, transmits mechanical effects of facial expressions to the ear region.
@@
By embedding an Inertial Measurement Unit (IMU) in an ear-mounted device, such as a wireless earbud, these changes can be detected. As facial expressions deform the ear canal, they alter the earbud’s position and orientation. The IMU captures these variations using accelerometers and gyroscopes, converting them into measurable signals. Each facial expression produces a distinct motion pattern, allowing the IMU to differentiate expressions based on the recorded data. Advanced machine learning algorithms process these signals to identify specific expressions, enabling unobtrusive and practical facial expression recognition. This approach capitalizes on the natural linkage between facial musculature and ear canal dynamics, offering a novel solution for wearable expression recognition systems.
4 Method
In Fig. 2, we present an overview of the architecture of our proposed IMUFace system, which is divided into two distinct phases: a training phase and a testing phase.
The training phase aims to teach a deep learning model to map IMU signals to facial movements using two synchronized data streams: video and IMU. The video stream provides ground truth by detecting and aligning faces in each frame, from which 2D facial landmarks are extracted. Meanwhile, the IMU stream, collected from earphone-embedded sensors, captures subtle facial muscle movements. This data undergoes a three-step preprocessing pipeline: a quick calibration to adapt to individual facial structures, a Short-Time Fourier Transform (STFT) to convert signals into the frequency domain, and windowed segmentation for temporal analysis. The processed IMU data is then fed into our transformer-based model, IMUTwinTrans, which predicts facial landmarks. A loss function measures the difference between the predicted and ground-truth landmarks, enabling the model to optimize its parameters through backpropagation.
Once training is complete, the system enters the testing phase for inference. In this phase, only the IMU data stream is required. The incoming IMU data undergoes the same preprocessing steps as in the training phase (calibration, STFT, and segmentation). This processed data is then fed into the fully trained model. The model, having learned the correlation between IMU signals and facial expressions, directly outputs the estimated facial landmarks in real-time. As the final step, these predicted landmarks are fitted to the parameters of the FLAME model[li2017learning] to reconstruct and animate a realistic 3D facial avatar.